# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ 4.1.3. Vulnerability Databases & Other Datasets
Another prominent source of software defects is pre-existing datasets, particularly those focused on security vulnerabilities, such as the Common Vulnerabilities and Exposures (CVE) dataset and the National Vulnerability Database (NVD). These datasets provide standardized, well-documented vulnerabilities, which are invaluable for creating benchmarks and conducting empirical studies. For example, CrossVul (Nikitopoulos et al., 2021) leverages CVE data to build comprehensive datasets of Java and cross-language vulnerabilities, aimed at studying program repair and vulnerability detection. Similarly, CVEjoin (da Ponte et al., 2023) consolidates information from CVE and other threat intelligence sources to create a unified dataset for security research. VulinOSS (Gkortzis et al., 2018) mines vulnerabilities from NVD and CVE to study security issues in open-source projects.
Besides the vulnerability databases mentioned above, some datasets rely on other well-established sources of defects. For example, both PreciseBugCollector (Ye et al., 2023a) and BugOss (Kim and Hong, 2023) build defect datasets from OSS-Fuzz (Serebryany, 2017), a continuous fuzzing platform by Google. These datasets, built upon existing structured vulnerability repositories, provide a solid foundation for evaluating and improving security practices, as well as for developing more robust vulnerability detection and repair techniques. However, these datasets are limited to previously identified defects, preventing the inclusion of newly emerging ones. Furthermore, their quality inherently depends on the underlying databases, which can vary significantly across different sources.
@@ 4.1.4. Developer Q&A Forums
Q&A platforms, e.g., Stack Exchange and Stack Overflow, provide valuable real-world insights into common bugs faced by developers. These platforms contain a wealth of developer-reported issues, which can be mined and categorized to capture bug symptoms, root causes, and their resolutions. For example, Morovati et al. (2023) and Chen et al. (2022) extract bugs from Q&A platforms to study defects in machine learning and blockchain systems, respectively. These datasets concentrate on practical, frequently encountered defects accompanied by detailed descriptions and resolutions, making them particularly valuable for empirical research. However, the unstructured nature of data on these platforms poses challenges for defect extraction, often requiring specialized expertise and manual processing efforts.
@@ 4.1.5. Coding Assignments
Another source of defects is submissions to coding assignments, programming competitions, and online judging platforms. While these defects may not reflect bugs in complex, real-world software, they often come with detailed descriptions and test cases, making them useful for evaluating bug detection and repair techniques. ConDefects (Wu et al., 2024) draws from coding platforms like AtCoder to collect bugs observed in a competitive programming environment. IntroClass (Le Goues et al., 2015) consists of bugs in student coding assignments and small C programs, providing benchmarks for automated program repair. Codeflaws (Tan et al., 2017) collects defects from competitive programming contests, while QuixBugs (Lin et al., 2017) is based on the Quixey programming competition, providing a multi-language benchmark for evaluating program repair techniques.
4.1.6. Continuous Integration & Delivery
A continuous integration and continuous delivery (CI/CD) system records histories of the outcome (e.g., failed or passed) of automated builds and tests after each code change, offering a valuable source of real-world failures and their resolutions. CI/CD is closely tied to version control systems, as it usually monitors code repositories to trigger builds and tests on each commit or pull request. By utilizing fail-pass patterns, several software defect datasets have been built to capture defects that arise during regular development cycles, ensuring their relevance for both research and practice. For instance, BugSwarm (Tomassi et al., 2019) and Bears (Madeiral et al., 2019) mine CI pipelines from Travis-CI, leveraging fail-pass transitions in the test outcomes to identify and record bug fixes. Similarly, with this model, BugSwarm (Tomassi et al., 2019; Zhu et al., 2023), GitBug-Java (Silva et al., 2024), and GitBug-Actions (Saavedra et al., 2024) also collect a wide range of defects across different software systems from GitHub Actions. These datasets emphasize reproducibility by providing detailed environments and scripts for setting up and replaying failures, making them valuable for execution-based benchmarking and debugging. Moreover, by utilizing the fail-pass pattern from CI/CD systems, these datasets capture real-world defects in an automated, reproducible manner, making them particularly valuable for studies that require large-scale, real-world, and reproducible defect datasets. Nevertheless, collecting defects from CI/CD pipelines inherently excludes defects that do not manifest within these automated processes, potentially overlooking issues that are caught before committing code into the repository or that manifest only in a production environment.
4.1.7. Synthesis
Some datasets generate defects through automated code mutation or bug injection. While this methodology cannot guarantee that the resulting defects are representative of real-world bugs, it provides a controlled and repeatable process for creating benchmarks for research. Specifically, it allows creators to control the location and types of defects, which is not feasible with naturally occurring bugs. Ngo et al. (2021) and RaceBench (Liang et al., 2023) introduce synthetic defects, such as variability and concurrency bugs, respectively, to evaluate bug localization and testing tools. FixReverter (Zhang et al., 2022a) focuses on bug injection by reverting CVE-related fixes to recreate real-world vulnerabilities for fuzz testing. Similarly, Bures et al. (2020) and Cai et al. (2024) inject faults into open-source projects and smart contracts, respectively, to assess security vulnerability detection. Additionally, with LLMs, DebugBench (Tian et al., 2024) injects faulty code into correct LeetCode solutions to create a defect dataset for evaluating the debugging ability of LLMs. These datasets provide a systematic way to study the effects of specific types of bugs and the efficacy of tools designed to detect or repair them.
@@ 4.2. Construction Methodology
In the following, we separate the construction process into three stages: defect collection, defect validation, and defect categorization. Defect collection involves gathering defects from the sources discussed in Section 4.1. Defect validation is the process of verifying that the collected defects genuinely represent software faults and defect categorization refers to assigning additional classification information to each defect within the dataset. Datasets with fully synthetic defects, as discussed in Section 4.1.7, are excluded from this section, since their construction does not involve defect mining methodologies. However, datasets that obtain code from external sources and subsequently mutate it, such as SBP (Cai et al., 2024), remain within the scope of this discussion.
@@ 4.2.1. Defect Collection
Keyword Matching (Timperley et al., 2024; Liu et al., 2024b; Jiang et al., 2024; Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Waseem et al., 2024; Das et al., 2024; Pramod et al., 2024; Song et al., 2023; Zhong et al., 2023; Avula et al., 2023; Mahbub et al., 2023; Morovati et al., 2023; Chen et al., 2023b; Guan et al., 2023; Xiong et al., 2023; Csuvik and Vidács, 2022; Kim et al., 2022; Zhang et al., 2022b; Paltenghi and Pradel, 2022; Wendland et al., 2021; Gyimesi et al., 2021; Yuan et al., 2021; Kim et al., 2021b; Paul et al., 2021; Kamienski et al., 2021; Makhshari and Mesbah, 2021; Sánchez et al., 2020; Zhang et al., 2020; Karampatsis and Sutton, 2020; Murphy et al., 2020; Bagherzadeh et al., 2020; Wang et al., 2020; Garcia et al., 2020; Benton et al., 2019; Radu and Nadi, 2019; Riganelli et al., 2019; Xu et al., 2019; Zhang et al., 2018; Liu et al., 2019; Lin et al., 2015; D’Ambros et al., 2012; Zaman et al., 2012; Chen et al., 2021b; Selakovic and Pradel, 2017; Xu et al., 2023; de Santana et al., 2024; Wang et al., 2023b; Azad et al., 2023; Wang et al., 2022b; Shen et al., 2021; Gu et al., 2019; Franco et al., 2017; Wang et al., 2017; Tan et al., 2018) 56
Mapping (Ruan et al., 2024; Ni et al., 2024; Jiang et al., 2024; Yu and Wang, 2024; Wang et al., 2024; Song et al., 2023; Zhong et al., 2023; Lee et al., 2024; Mahbub et al., 2023; Chen et al., 2023b; Jiang et al., 2023; Bhuiyan et al., 2023; Guan et al., 2023; Xiong et al., 2023; Liang et al., 2022; Keshavarz and Nagappan, 2022; Bui et al., 2022; Pereira et al., 2022; Akimova et al., 2021; Kim et al., 2021b; Paul et al., 2021; Ferenc et al., 2020; Wu et al., 2020; Vieira et al., 2019; Riganelli et al., 2019; Ponta et al., 2019; Lin et al., 2019; Saha et al., 2018; Gkortzis et al., 2018; Lin et al., 2015; Just et al., 2014; D’Ambros et al., 2012; Jalbert et al., 2011; Ho et al., 2023; Wang et al., 2023a, c, 2021a; Ren et al., 2020; Franco et al., 2017; Wang et al., 2017; Tan et al., 2018) 41
Manual Collection (Timperley et al., 2024; Liu et al., 2024b; Zhao et al., 2023a; Applis and Panichella, 2023; Liu et al., 2022; Ponta et al., 2019; Lin et al., 2019; Ohira et al., 2015; Jalbert et al., 2011; Ku et al., 2007; Gunawi et al., 2014) 11
External Oracle (Liu et al., 2024a; Tian et al., 2024; Silva et al., 2024; Wu et al., 2024; Kim and Hong, 2023; Saavedra et al., 2024; Tomassi et al., 2019; Madeiral et al., 2019; Tan et al., 2017; Le Goues et al., 2015) 10
SZZ Algorithm (Śliwerski et al., 2005) (Avula et al., 2023; Mahbub et al., 2023; Keshavarz and Nagappan, 2022; Paul et al., 2021; Ferenc et al., 2020; Karampatsis and Sutton, 2020; Ren et al., 2020) 7
Static Analysis (Wang et al., 2024; Senanayake et al., 2023; Pereira et al., 2022; Zheng et al., 2021; Ferenc et al., 2020; Karampatsis and Sutton, 2020) 6
Machine Learning (Wang et al., 2024; Zheng et al., 2021; Wu et al., 2020) 3
Others (Zhang et al., 2024a; Jimenez et al., 2024; Cai et al., 2024; Brust et al., 2023; da Ponte et al., 2023; Chen et al., 2022; Song et al., 2022; Afrose et al., 2023; Du et al., 2021; Widyasari et al., 2020; Gao et al., 2018; Madeyski and Kawalerowicz, 2017; Le Goues et al., 2015; Lamkanfi et al., 2013; Zhao et al., 2023b; Eghbali and Pradel, 2020) 16
@@
Defect collection refers to the process of gathering defects from one or more mining sources. The particular strategies that dataset creators employ depend on their choice of defect sources and on the intended scope of the dataset. For instance, if a dataset is sourced from CI/CD systems, then collecting defects based on historical test failures is a natural choice. Table 6 lists several common defect collection strategies, showing the diversity of approaches used to construct software defect datasets. Keyword matching and mapping are the most common strategies, with 56 and 41 datasets using them, respectively. Other strategies, such as manual collection, external oracle use, and specialized algorithms (Śliwerski et al., 2005), are less common, but still used by a number of datasets.
@@
Many datasets attempt to find defects by scanning for specific keywords (e.g., “bug” or “fix”) in commit messages, issue reports, and other relevant text fields. Zhong et al. (2023) make for a prototypical example: they examine all GitHub push and pull request events in their target repositories, and consider a bug-fixing event to be any commit or pull request that contains one of the terms “fix,” “repair,” “patch,” “bug,” “issue,” “problem,” or “fault” in the commit message or pull request description. Those keywords are all typical for this kind of approach, built off the assumption that a commit that mentions bugs, repairs, and so on is likely attempting to fix a defect. Keyword-based defect collection has limited precision (Tian et al., 2012), but it is still one of the most common collection approaches observed in this survey, presumably due to its simplicity and its abundance (56 datasets, 56/132) in previous literature.
@@
Another common approach in defect collection is to map defects from one source to objects from another source, such as mapping commits from a project’s Git history to related bug reports documented in its issue tracker. This strategy is necessary for datasets that (1) source defects from bug reports or vulnerability databases (Sections 4.1.1 and 4.1.3) and (2) wish to include source-code-level information (Section 5.2). Accordingly, mapping is common, with 41 datasets (31.1%) utilizing it. The difficulty of this process can vary based on the mapping source and destination. For instance, mapping GitHub issues to GitHub commits is quite easy, since GitHub tracks commits that reference issues and exposes that relationship in its website and API. In contrast, mapping CVE entries to fixing commits can be more challenging, since many CVE datasets do not reference the commits that fix a vulnerability, and likewise many vulnerability-fixing commits do not reference their associated CVEs. Jiang et al. (2024) outline their process for mapping CVEs to bug-fixing commits in detail, using six different strategies to construct these relationships.
@@
Some dataset creators forego automatic defect collection, but instead manually collect defects. For instance, Jalbert et al. (2011) create RADBench by searching through the issue trackers of several projects to find concurrency bugs. An advantage of this approach is that it allows researchers to make qualitative decisions about what defects to include in their datasets, such as only including “vital” issues, which can depend on a variety of factors that are not easy to judge automatically. Conversely, a notable disadvantage is the substantial increase in manual effort imposed on dataset creators. Consequently, it is unsurprising that manually collected datasets typically exhibit a smaller median size compared to automatically collected datasets (298 versus 1158 entries, respectively). Also, while qualitative judgments about a defect can be useful, they are also inherently subjective, which may deter potential dataset users who prioritize objectivity. These disadvantages are probably why this strategy is, with only eleven datasets, relatively uncommon among the surveyed datasets.
@@
Some datasets use historical test failures or continuous integration (CI) failures as external oracles to distinguish buggy from non-buggy software. This should not be confused with datasets that reproduce test failures, which are described in Section 4.2.2. For example, BugSwarm (Tomassi et al., 2019; Zhu et al., 2023) is created by scanning through the commit history of repositories that use GitHub Actions or Travis-CI, looking for cases where one commit’s CI run fails and a subsequent commit’s run passes. Using historical failures to distinguish bugs from non-bugs allows for better confidence in a dataset’s accuracy, given that failing tests are a tell-tale sign of a bug. However, since this strategy relies on historical test or CI results, it inherently limits the bugs that can appear in a dataset to those that have this historical information available. This naturally limits the defect source of a dataset, which can be an issue if the dataset creators are concerned about dataset size. Moreover, in the case of CI failures, care must be taken to ensure that a given failure was actually caused by a failing test (as opposed to, say, a code style checker).
@@
The SZZ algorithm, introduced by Śliwerski et al. (2005) and refined multiple times since (Kim et al., 2006; Williams and Spacco, 2008), attempts to identify the software change that introduced a bug given the change that fixes the bug. This addresses a weakness in some other collection strategies (e.g., keyword search), which is that while they can readily find software changes that fix bugs, they have no way to identify the change(s) that caused those bugs. There are seven datasets in our survey using SZZ or a similar algorithm to mine bug-introducing commits. Mahbub et al. (2023) demonstrate a typical approach: after finding bug-fixing pull requests in their target repositories, they use SZZ to associate one or more bug-introducing commits to each bug fix.

### v2

@@ 4.1.3. Vulnerability Databases & Other Datasets
Another prominent source of software defects is pre-existing datasets, particularly those focused on security vulnerabilities, such as the Common Vulnerabilities and Exposures (CVE) dataset and the National Vulnerability Database (NVD). These datasets provide standardized, well-documented vulnerabilities, which are invaluable for creating benchmarks and conducting empirical studies. For example, CrossVul (DBLP:conf/sigsoft/NikitopoulosDLM21) leverages CVE data to build comprehensive datasets of Java and cross-language vulnerabilities, aimed at studying program repair and vulnerability detection. Similarly, CVEjoin (DBLP:conf/aina/PonteRM23a) consolidates information from CVE and other threat intelligence sources to create a unified dataset for security research. VulinOSS (DBLP:conf/msr/GkortzisMS18) mines vulnerabilities from NVD and CVE to study security issues in open-source projects.
Besides the vulnerability databases mentioned above, some datasets rely on other well-established sources of defects. For example, both PreciseBugCollector (DBLP:conf/kbse/YeCG23) and BugOss (DBLP:conf/icst/KimH23) build defect datasets from OSS-Fuzz (Serebryany, 2017), a continuous fuzzing platform by Google. These datasets, built upon existing structured vulnerability repositories, provide a solid foundation for evaluating and improving security practices, as well as for developing more robust vulnerability detection and repair techniques. However, these datasets are limited to previously mined defects, preventing the inclusion of newly emerging ones. Furthermore, their quality inherently depends on the underlying databases, which can vary significantly across different sources.
@@ 4.1.4. Developer Q&A Forums
Q&A platforms, e.g., Stack Exchange and Stack Overflow, provide valuable real-world insights into common bugs faced by developers. These platforms contain a wealth of developer-reported issues, which can be mined and categorized to capture bug symptoms, root causes, and their resolutions. For example, DBLP:journals/ese/MorovatiNKJ23 and DBLP:journals/tse/ChenXLGLC22 extract bugs from Q&A platforms to study defects in machine learning and blockchain systems, respectively. These datasets concentrate on practical, frequently encountered defects accompanied by detailed descriptions and resolutions, making them particularly valuable for empirical research. However, the unstructured nature of data on these platforms poses challenges related to data quality (Ponzanelli et al., 2014), readability (Duijn et al., 2015), and human-factors (Wang et al., 2020), while LLM-based approaches show promise in automating the extraction and structuring of defects from such sources (da Silva et al., 2025).
4.1.5. Continuous Integration & Delivery
A continuous integration and continuous delivery (CI/CD) system records histories of the outcome (e.g., failed or passed) of automated builds and tests after each code change, offering a valuable source of real-world failures and their resolutions. CI/CD is closely tied to version control systems, as it usually monitors code repositories to trigger builds and tests on each commit or pull request. By utilizing fail-pass patterns, several software defect datasets have been built to capture defects that arise during regular development cycles, ensuring their relevance for both research and practice. For instance, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19) and Bears (DBLP:conf/wcre/DelfimUMM19) mine CI pipelines from Travis-CI, leveraging fail-pass transitions in the test outcomes to identify and record bug fixes. Similarly, with this model, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19; Zhu et al., 2023), GitBug-Java (DBLP:conf/msr/SilvaSM24), and GitBug-Actions (DBLP:conf/icse/SaavedraSM24) also collect a wide range of defects across different software systems from GitHub Actions. These datasets emphasize reproducibility by providing detailed environments and scripts for setting up and replaying failures, making them valuable for execution-based benchmarking and debugging. Moreover, by utilizing the fail-pass pattern from CI/CD systems, these datasets capture real-world defects in an automated, reproducible manner, making them particularly valuable for studies that require large-scale, real-world, and reproducible defect datasets. Nevertheless, collecting defects from CI/CD pipelines inherently excludes defects that do not manifest within these automated processes, potentially overlooking issues that are caught before committing code into the repository or that manifest only in a production environment.
4.1.6. Synthesis
Some datasets generate defects through automated code mutation or bug injection. While this methodology cannot guarantee that the resulting defects are representative of real-world bugs, it provides a controlled and repeatable process for creating benchmarks for research. Specifically, it allows creators to control the location and types of defects, which is not feasible with naturally occurring bugs. DBLP:conf/splc/NgoNNV21 and RaceBench (DBLP:conf/asiaccs/LiangYDMHZ23) introduce synthetic defects, such as variability and concurrency bugs, respectively, to evaluate bug localization and testing tools. FixReverter (DBLP:conf/uss/ZhangP0W22) focuses on bug injection by reverting CVE-related fixes to recreate real-world vulnerabilities for fuzz testing. Similarly, DBLP:conf/icst/BuresHA20 and DBLP:journals/jss/CaiLZZS24 inject faults into open-source projects and smart contracts, respectively, to assess security vulnerability detection. Additionally, with LLMs, DebugBench (DBLP:conf/acl/TianYQCLPWHL0024) injects faulty code into correct LeetCode solutions to create a defect dataset for evaluating the debugging ability of LLMs. These datasets provide a systematic way to study the effects of specific types of bugs and the efficacy of tools designed to detect or repair them.
@@ 4.1.7. Coding Assignments
Another source of defects is submissions to coding assignments, programming competitions, and online judging platforms. While these defects may not reflect bugs in complex, real-world software, they often come with detailed descriptions and test cases, making them useful for evaluating bug detection and repair techniques. ConDefects (DBLP:conf/sigsoft/WuLZ024) draws from coding platforms like AtCoder to collect bugs observed in a competitive programming environment. IntroClass (DBLP:journals/tse/GouesHSBDFW15) consists of bugs in student coding assignments and small C programs, providing benchmarks for automated program repair. Codeflaws (DBLP:conf/icse/TanYYMR17) collects defects from competitive programming contests, while QuixBugs (DBLP:conf/oopsla/LinKCS17) is based on the Quixey programming competition, providing a multi-language benchmark for evaluating program repair techniques.
@@ 4.2. Construction Methodology
In the following, we separate the construction process into three stages: defect collection, defect validation, and defect categorization. Defect collection involves gathering defects from the sources discussed in Section 4.1. Defect validation is the process of verifying that the collected defects genuinely represent software faults and defect categorization refers to assigning additional classification information to each defect within the dataset. Datasets with fully synthetic defects, as discussed in Section 4.1.6, are excluded from this section, since their construction does not involve defect mining methodologies. However, datasets that obtain code from external sources and subsequently mutate it, such as SBP (DBLP:journals/jss/CaiLZZS24), remain within the scope of this discussion.
@@ 4.2.1. Defect Collection
Keyword Matching (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/raid/0001DACW23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/CsuvikV22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/icse/Makhshari021; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/hotsos/MurphyBSR20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/dsa/XuYWA19; DBLP:conf/issta/ZhangCCXZ18; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/msr/ZamanAH12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23) 64
Mapping (DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/qrs/YuW24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icst/LeeKYY24; DBLP:conf/msr/MahbubSR23; DBLP:conf/raid/0001DACW23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/BuiSF22; DBLP:conf/prdc/PereiraAV22; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:journals/jss/FerencGGTG20; DBLP:journals/jss/WuZCWM20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/msr/GkortzisMS18; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/issta/JustJE14; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/iwpc/RenL020; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/issta/WuJPLD0BS23) 42
Manual Collection (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/PontaPSBD19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/kbse/KuHCL07; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/OliverDAH24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/tse/ZhaoXBCL23) 16
External Oracle (DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/msr/SilvaSM24; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/icst/KimH23; DBLP:conf/icse/SaavedraSM24; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/kbse/XuZL23) 13
SZZ Algorithm (Sliwerski et al., 2005) (DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:conf/msr/KeshavarzN22; DBLP:conf/icse/PaulTB21a; DBLP:journals/jss/FerencGGTG20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/iwpc/RenL020) 7
Static Analysis (DBLP:conf/icse/WangHGWC024; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/prdc/PereiraAV22; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:journals/jss/FerencGGTG20; DBLP:conf/msr/KarampatsisS20) 6
Machine Learning (DBLP:conf/icse/WangHGWC024; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:journals/jss/WuZCWM20) 3
Others (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:journals/jss/CaiLZZS24; DBLP:journals/compsec/BrustSG23; DBLP:conf/aina/PonteRM23a; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/issre/Du0MZ21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/msr/MadeyskiK17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/LamkanfiPD13; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/kbse/EghbaliP20; DBLP:conf/msr/BeyerGKRR24) 17
@@
Defect collection refers to the process of gathering defects from one or more mining sources. The particular strategies that dataset creators employ depend on their choice of defect sources and on the intended scope of the dataset. For instance, if a dataset is sourced from CI/CD systems, then collecting defects based on historical test failures is a natural choice. Table 7 lists several common defect collection strategies, showing the diversity of approaches used to construct software defect datasets. Keyword matching and mapping are the most common strategies, with 64 and 42 datasets using them, respectively. Other strategies, such as manual collection, external oracle use, and specialized algorithms (Sliwerski et al., 2005), are less common, but still used by a number of datasets.
@@
Many datasets attempt to find defects by scanning for specific keywords (e.g., “bug” or “fix”) in commit messages, issue reports, and other relevant text fields. DBLP:conf/kbse/ZhongLZGWGL23 make for a prototypical example: they examine all GitHub push and pull request events in their target repositories, and consider a bug-fixing event to be any commit or pull request that contains one of the terms “fix,” “repair,” “patch,” “bug,” “issue,” “problem,” or “fault” in the commit message or pull request description. Those keywords are all typical for this kind of approach, built off the assumption that a commit that mentions bugs, repairs, and so on is likely attempting to fix a defect. Keyword-based defect collection has limited precision (Tian et al., 2012), but it is still one of the most common collection approaches observed in this survey, presumably due to its simplicity and its abundance (64 datasets, 64/151) in previous literature.
@@
Another common approach in defect collection is to map defects from one source to objects from another source, such as mapping commits from a project’s Git history to related bug reports documented in its issue tracker. This strategy is necessary for datasets that (1) source defects from bug reports or vulnerability databases (Sections 4.1.1 and 4.1.3) and (2) wish to include source-code-level information (Section 5.2). Accordingly, mapping is common, with 42 datasets (42/151) utilizing it. The difficulty of this process can vary based on the mapping source and destination. For instance, mapping GitHub issues to GitHub commits is quite easy, since GitHub tracks commits that reference issues and exposes that relationship in its website and API. In contrast, mapping CVE entries to fixing commits can be more challenging, since many CVE datasets do not reference the commits that fix a vulnerability, and likewise many vulnerability-fixing commits do not reference their associated CVEs. DBLP:journals/tosem/JiangJWMLZ24 outline their process for mapping CVEs to bug-fixing commits in detail, using six different strategies to construct these relationships.
@@
Some dataset creators forego automatic defect collection, but instead manually collect defects. For instance, DBLP:conf/hotpar/JalbertPPS11 create RADBench by searching through the issue trackers of several projects to find concurrency bugs. An advantage of this approach is that it allows researchers to make qualitative decisions about what defects to include in their datasets, such as only including “vital” issues, which can depend on a variety of factors that are not easy to judge automatically. Conversely, a notable disadvantage is the substantial increase in manual effort imposed on dataset creators. Consequently, it is unsurprising that manually collected datasets typically exhibit a smaller median size compared to automatically collected datasets (298 versus 1158 entries, respectively). Also, while qualitative judgments about a defect can be useful, they are also inherently subjective, which may deter potential dataset users who prioritize objectivity. These disadvantages are probably why this strategy is, with only 16 datasets, relatively uncommon among the surveyed datasets.
@@
Some datasets use historical test failures or continuous integration (CI) failures as external oracles to distinguish buggy from non-buggy software. This should not be confused with datasets that reproduce test failures, which are described in Section 4.2.2. For example, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19; Zhu et al., 2023) is created by scanning through the commit history of repositories that use GitHub Actions or Travis-CI, looking for cases where one commit’s CI run fails and a subsequent commit’s run passes. Using historical failures to distinguish bugs from non-bugs allows for better confidence in a dataset’s accuracy, given that failing tests are a tell-tale sign of a bug. However, since this strategy relies on historical test or CI results, it inherently limits the bugs that can appear in a dataset to those that have this historical information available. This naturally limits the defect source of a dataset, which can be an issue if the dataset creators are concerned about dataset size. Moreover, in the case of CI failures, care must be taken to ensure that a given failure was actually caused by a failing test (as opposed to, say, a code style checker).
@@
The SZZ algorithm, introduced by Sliwerski et al. (2005) and refined multiple times since (Kim et al., 2006; DBLP:conf/issta/WilliamsS08), attempts to identify the software change that introduced a bug given the change that fixes the bug. This addresses a weakness in some other collection strategies (e.g., keyword search), which is that while they can readily find software changes that fix bugs, they have no way to identify the change(s) that caused those bugs. There are seven datasets in our survey using SZZ or a similar algorithm to mine bug-introducing commits. DBLP:conf/msr/MahbubSR23 demonstrate a typical approach: after finding bug-fixing pull requests in their target repositories, they use SZZ to associate one or more bug-introducing commits to each bug fix.